# ZMM-TTS: Zero-shot Multilingual and Multispeaker Speech Synthesis Conditioned on Self-supervised Discrete Speech Representations

### v1

@@ VI-B Experimental results and analysis
The results on unseen languages are shown in Table VII.
@@ VI-B1 Comparison between different input representations
In high-resource baselines, the CER/WER results of ZMM-TTS2x and ZMM-TTS2c were quite similar. However, when only 15 minutes of data were used for fine-tuning, the synthesized speech produced by ZMM-TTS2x was much more intelligible than that of ZMM-TTS2c in low-resource scenarios. The same phenomenon is also observed in the case of FSM2x and FSM2c. This result demonstrates that the pre-trained phoneme representations from XPhoneBert are more suitable for low-resource scenarios.
@@ VI-B2 Comparison of FSM and ZMM-TTS models
The model FSM2x outperforms ZMM-TTS2x in terms of CER/WER. This observation is consistent across six high-resource languages as well, as discussed in Section V-B. However, in terms of AMOS and SECS, the ZMM-TTS2x model produces significantly better speech naturalness and speaker similarity than the FSM2x model, both in high and low resource scenarios on two unseen languages. This result indicates that self-supervised features can better reconstruct sound quality and timbre compared to the Mel spectrum, even with limited data.
@@ VI-B3 Comparison of ZMM-TTS and YourTTS models
We can see that YourTTS does not have the ability to be finetuned with limited data from an unseen language. Even if one hour of data is used for finetuning, YourTTS still cannot synthesize understandable audio. Although promising results are evident in some language combinations with sufficient training data, YourTTS demonstrates limited efficacy when adapting to languages with limited available speakers.
@@ VI-B4 Comparison of different data size
As more training data is added, the synthesized speech quality improves, particularly in terms of speech intelligibility. Furthermore, the results for speaker similarity SECS of ZMM-TTS2x are similar across different quantities of training data (2.5 min, 5 min and 15 min). In general, with only 2.5 minutes of speech from an unseen speaker in a new language, our fine-tuning of the ZMM-TTS2x still resulted in high speaker similarity.
@@ VI-B5 Comparison of two languages
We found that for Italian, using pre-trained phoneme representations as input, we were able to synthesize intelligible speech in zero-shot scenarios. In Polish, there is a significant gap between the zero-shot and few-shot scenarios, but both FSM2x and ZMM-TTS2x were able to generate intelligible speech utilizing a few minutes of data with pre-trained phoneme representations as input. The SSL speech-based features do not seem to provide an advantage over the Mel spectrograms, unlike in Italian. Although Italian and Polish are both included in the XLSR training data, it is possible that the performance of SSL representations is heavily influenced by the similarity between the domains of the ZMM-TTS pre-training data and the fine-tuning data. Future work may include investigating the impact of language similarity on cross-lingual transfer.
@@ VII Conclusion
In this paper, we propose a new method for generating multilingual speech using self-supervised discrete speech representations. We experimented with different input representations and incorporated a pre-trained multilingual phoneme encoder for multilingual tasks. Our experimental results show that our approach improves the speaker similarity and naturalness of synthetic speech in multilingual tasks, even for unseen speakers. Additionally, our framework achieves high intelligibility and speaker similarity in limited-data or zero-shot scenarios for a new language. In the future, we plan to apply this model to more languages and try advanced language and speaker adaptation strategies.

### v2

@@ VI-B Experimental results and analysis
The results on unseen languages are shown in Table VIII.
@@ VI-B1 Comparison between different input representations
In high-resource baselines, the intelligibility test results (CER/IMOS) of ZMM-TTS2x and ZMM-TTS2c were quite similar. However, when only 15 minutes of data were used for fine-tuning, the synthesized speech produced by ZMM-TTS2x was much more intelligible than that of ZMM-TTS2c in low-resource scenarios. The same phenomenon is also observed for FSM2x and FSM2c. This result demonstrates that the pre-trained phoneme representations from XPhoneBert are more suitable for low-resource scenarios.
@@ VI-B2 Comparison of FSM and ZMM-TTS models
On both subjective and objective evaluation, the ZMM-TTS2x model produces significantly more natural speech (MOS/UTMOS) with better speaker similarity (DMOS/SECS) than the FSM2x model, both in high and low resource scenarios on two unseen languages. This result indicates that self-supervised features can better reconstruct sound quality and timbre compared with the Mel spectrograms, even with limited data. Furthermore, on the basis of the subjective IMOS evaluations, in low-resource scenarios, the intelligibility of ZMM-TTS2x surpasses that of FSM2x. This demonstrates the advantage of our model in low-resource scenarios.
@@ VI-B3 Comparison of ZMM-TTS and YourTTS models
We can see that YourTTS cannot be fine-tuned with limited data from an unseen language. Even if one hour of data is used for fine-tuning, YourTTS still cannot synthesize understandable audio. Although promising results are evident in several language combinations with sufficient training data, YourTTS demonstrates limited efficacy when adapting to languages with limited available speakers.
@@ VI-B4 Comparison of different data size
As more training data is added, the synthesized speech quality improves, particularly in terms of speech intelligibility. Furthermore, the results for speaker similarity (SECS/DMOS) of ZMM-TTS2x are similar across different quantities of training data (2.5, 5, and 15 mins). In general, with only 2.5 mins of speech from an unseen speaker in a new language, our fine-tuning of the ZMM-TTS2x still resulted in high speaker similarity in terms of SECS.
@@ VI-B5 Comparison of two languages
We found that for Italian, using pre-trained phoneme representations as input, we were able to synthesize intelligible speech in zero-shot scenarios. In Polish, there is a significant gap between the zero-shot and few-shot scenarios, but both FSM2x and ZMM-TTS2x were able to generate intelligible speech utilizing a few minutes of data with pre-trained phoneme representations as input. The SSL speech-based features do not seem to provide an advantage over the Mel spectrograms, unlike in Italian. Although Italian and Polish are both included in the XLSR training data, the performance of SSL representations may be heavily affected by the similarity between the domains of the ZMM-TTS pre-training data and the fine-tuning data. Future work may include investigating the impact of language similarity on cross-lingual transfer.
@@ VII Conclusion
In this paper, we propose a new method for generating multilingual speech using self-supervised discrete speech representations. We experimented with different input representations and incorporated a pre-trained multilingual phoneme encoder for multilingual tasks. Our experimental results show that our approach improves the speaker similarity and naturalness of synthetic speech in multilingual tasks, even for unseen speakers. Additionally, our framework achieves high intelligibility and speaker similarity in limited-data or zero-shot scenarios for a new language.
As future work, we plan to apply this model to many more languages and explore advanced language and speaker adaptation strategies. To make this happen, we will also need to investigate flexible language representations to handle unseen languages more accurately than the language IDs used in the current system.
Appendix A Comparison with large-scale TTS systems
Considering the remarkable breakthroughs achieved by large-scale TTS models for monolingual speech synthesis, despite our model primarily targeting multilingual and low-resource scenarios, we still conducted comparative analyses from various perspectives on the English language, in contrast to these models.
A-A Experimental conditions
A-A1 Implementation details
For a fair comparison, in addition to the ZMM-TTS2x model trained on six languages in Section IV, we also trained another model on the LibriTTS [85] dataset. We compared our ZMM-TTS2x with the following three strong zero-shot TTS baselines:
  • VALL-E-X[66]. It is a multilingual version of VALL-E that also employs a combination of autoregressive and non-autoregressive methods for discrete token generation. We use an open-source implementation and checkpoint  [^10] https://github.com/Plachtaa/VALL-E-X. The training data for this model comprises three languages: Chinese (ch), English (en), and Japanese (jp).
  • HierSpeech++[32]. This work is an extended version of HierSpeech [64]. It uses a non-autoregressive model for continuous vector generation. There are three components in HierSpeech++: a hierarchical speech synthesizer, text-to-vec (TTV), and speech super-resolution (SpeechSR). We conducted experiments using two different pre-trained HierSpeech++ models, which were labeled as HierSpeech++(a) and HierSpeech++(b). The difference between them lies in the training data scale of the hierarchical speech synthesizer. HierSpeech++(a) used the complete LibriTTS train-clean data, while HierSpeech++(b) utilized LibriTTS train-clean-100 and 360 subsets. However, the training data for TTV in both HierSpeech++(a) and HierSpeech++(b) remained consistent, employing the complete LibriTTS train-clean data. Note that for purposes of fair comparison, we did not use the super-resolution model. We used the official code and checkpoint for the experiments  [^11] https://github.com/sh-lee-prml/HierSpeechpp.
  • StyleTTS 2[65]. It leverages style diffusion and adversarial training with large speech language models (12-layer WavLM pre-trained on 94k hours of data) to achieve human-level TTS synthesis. We use the official code and checkpoint  [^12] https://github.com/yl4579/StyleTTS2.
A-A2 Evaluation dataset
We chose LibriSpeech [86] test-clean as our benchmark dataset for the zero-shot TTS task. This widely-used test set comprises 40 different speakers and 5.4 hours of speech. To conduct the benchmark experiments, we followed the approach described in [31], and we randomly selected 25 sentences for each speaker from the LibriSpeech test-clean dataset.
A-A3 Metrics
We evaluate the intelligibility of synthesized speech using WER as described in Section IV-D. Following previous works [22, 31], we compute SECS using the SOTA speaker verification model, WavLM-Large [^13] https://github.com/microsoft/UniSpeech/tree/main/
downstreams/speaker_verification
, to evaluate the speaker similarity, enabling comparison with those studies. We also use UTMOS like Section VI. In addition to evaluating the speech quality, we also measure the efficiency of the proposed model based on its Real-Time Factor (RTF), which is the time it takes to generate one second of audio on GPU and the number of parameters used in the model. We test the RTF on a single GPU (NVIDIA RTX 4090 with 24GB memory). Note that VALL-E-X and StyleTTS 2 generate audio using a sample rate of 24 kHz, while ZMM-TTS2x and HierSpeech++ use a sample rate of 16 kHz. To ensure consistency across all audio, we resampled it to 16 kHz and applied amplitude normalization using sv56 before conducting the WER, SECS, and UTMOS tests.
TABLE IX: Evaluation results for ZMM-TTS and recent large-scale TTS systems on LibriSpeech test-clean. LT-460 denotes LibriTTS train-clean-100 and 360 subsets, and LT-960 additionally utilizes a train-other-500 subset with LT-460. AL-1&3 denotes AISHELL-1 and AISHELL-3, and JP-CV denotes the Japanese Common Voice.
Method Training corpus Evaluation metric
Datasets Language Speaker Hour WER SECS UTMOS Params RTF
VALL-E-X LT-960, AL-1&3, JP-CV, others en, jp, ch $>3,122$ $1,739$ 26.77 0.512 3.29 395M 5.917
HierSpeech++(a) LT-960 en 2,311 555 2.03 0.591 4.40 204M 0.217
HierSpeech++(b) LT-460 en 1,151 245 2.17 0.555 4.38 204M 0.222
StyleTTS 2 LT-460 en 1,151 245 3.06 0.455 4.23 191M 0.070
ZMM-TTS2x(a) MLS, GLB, LJ, CSS10, NST en, fr, ge, pt, es, sw 546 130 4.06 0.432 3.89 167M 0.003
ZMM-TTS2x(b) LT-960 en 2,311 555 2.37 0.644 4.07 167M 0.003
Ground-truth - - - - 2.14 - 4.13 - -
A-B Experimental results and analysis
Table IX contains information on the efficiency and quality of all models being compared.
A-B1 SECS analysis
We found that models VALL-E-X, HierSpeech++(a), and HierSpeech++(b) outperform our model ZMM-TTS2x (a) in terms of unseen speaker similarity on LibriSpeech test sets. This gap may be due to differences in the scale of training data. ZMM-TTS2x (a) training corpus includes only 546 speakers, while other baseline systems have over 1,000 or 2,000 speakers in their training data. When reducing the scale of training data, the SECS of HierSpeech++(b) has a significant decline compared with HierSpeech++(a). With the same data scale, our proposed ZMM-TTS2x (b) achieved the best speaker similarity SECS compared with other recent zero-shot models. This demonstrates that our proposed model, which utilizes a larger amount of data, scales in performance and can generalize to unseen speakers.
A-B2 UTMOS analysis
After analyzing the scores predicted by the automatic MOS evaluation model, we noticed a correspondence between the quality of generated audio and the training corpus. Models trained on a clean TTS database LT-960 or LT-640 all achieved better UTMOS values. The training corpora for both the ZMM-TTS2x (a) model and VALL-E-X model do not exclusively consist of high-quality clean audio. They also contain some data that includes noticeable noise, such as the GlobalPhone in ZMM-TTS2x (a) and other databases in VALL-E-X.
A-B3 CER analysis
We have observed that the speech intelligibility achieved by the VALL-E-X model is the worst among all the models. This is primarily due to the autoregressive process of the model, which often results in synthesized speech containing errors such as skipped words and repeated words. On the other hand, our ZMM-TTS2x (a) model, despite being trained on a smaller corpus and multiple languages, has achieved WER results in English that are comparable with other models. This indicates the robustness of our model.
A-B4 Latency analysis
Undoubtedly, the autoregressive process of VALL-E-X results in the slowest synthesis speed, making real-time synthesis challenging to achieve on GPUs. Due to its non-autoregressive structure, ZMM-TTS2x is capable of real-time high-speed synthesis on GPU. Although HierSpeech++ is also non-autoregressive, the iterative process of the diffusion model still impacts the synthesis speed. Additionally, ZMM-TTS2x has around 167M parameters, thus being smaller than other TTS systems we compare.
Note that the baseline models focus solely on resource-rich languages such as English. In contrast, our model has strong language adaptability, making it more suitable for low-resource languages.